OpenStudy (anonymous):

Find the slope of the line through the pair of points. (6,-9) (-3,5)

OpenStudy (anonymous):

\[\frac{-9-5}{6+3}=-\frac{14}{9} \]

OpenStudy (anonymous):

Plot attached.